OpenStudy (anonymous):

Integrate sqrt{5sin^{2}(2t)} from 0 to pi

OpenStudy (watchmath):

\(\sin^2(t)=\frac{1}{2}(1-cos(2t))\) So \(\int 5\sin^2(2t)\,dt=\int \frac{5}{2}(1-\cos(4t))\,dt=\frac{5}{2}t-\frac{5}{8}\sin(4t)+C\)

OpenStudy (watchmath):

Oh sorry this is a definite integral... so just plugin the limits of integration :).

OpenStudy (anonymous):

How did you eliminate the square root?

OpenStudy (watchmath):

I am sorry, I didn't see the square root ... :)

OpenStudy (watchmath):

Ok, then it is simply the integral \[\sqrt{5}\int_0^\Pi \sin(2t)\,dt\] I believe you can do this integral now :).

OpenStudy (watchmath):

Sorry ... I made a mistake

OpenStudy (watchmath):

Over the interval \([0,\pi]\) the function \(\sin(2t\) is above the \(t-axis\) on \([0,\pi/2]\) and below the \(t\)-axis on \([\pi/2,\pi]\). So \(\sqrt{5\sin^2{2t}}=\sqrt{5}\sin{2t}\) on \([0,\pi/2]\) and equal to \(-\sqrt{5}\sin(2t)\) on the interval \([\pi/2,\pi]\). So the integral si equal to \[ \int_0^{\pi/2} \sqrt{5}\sin(2t)\,dt-\int_{\pi/2}^{\pi} \sqrt{5}\sin(2t)\,dt\]

OpenStudy (anonymous):

Thank you so much, this is the part I was missing.